OpenStudy (anonymous):

A hexagon is inscribed in a circle. If the difference between the area of the circle and the area of the hexagon is 36 m2, use the formula for the area of a sector to approximate the radius r of the circle. (Round your answer to three decimal places.)

OpenStudy (anonymous):

Acircle - Ahexagon = pi*r² - 6*r*(r*sqrt{3})/4)/2 = 36

OpenStudy (raden):

that should be ........... sqrt{3})/2)/2 = 36, right @thearcane ?

OpenStudy (anonymous):

Yep, I wrote wrong. Thanks.